On March 1, 2021, Gold Examiner receives $150,000 from a local bank and promises to deliver 95 units of certified 1-oz. gold bars on a future date. The contract states that ownership passes to the bank when Gold Examiner delivers the products to Brink’s, a third-party carrier. In addition, Gold Examiner has agreed to provide a replacement shipment at no additional cost if the product is lost in transit. The stand-alone price of a gold bar is $1,520 per unit, and Gold Examiner estimates the stand-alone price of the replacement insurance service to be $80 per unit. Brink’s picked up the gold bars from Gold Examiner on March 30, and delivery to the bank occurred on April 1.

Required:
1. How many performance obligations are in this contract?
2. to 4. Prepare the journal entry Gold Examiner would record on March 1, March 30 and April 1.
Answer #1
1
Number of performance obligations in the contract is 2.
2 to 4
Debit Credit
March 01, 2021 Cash 150000
     Deferred revenue—gold bars 142500
     Deferred revenue—insurance 7500
March 30, 2021 Deferred revenue—gold bars 142500
        Sales revenue 142500
April 01, 2021 Deferred revenue—insurance 7500
       Service revenue 7500
Workings:
Value of the gold bars 144400 =1520*95
Stand­alone selling price of the insurance 7600 =80*95
Total of stand­alone prices 152000
Allocation:
Deferred revenue—gold bars 142500 =150000*144400/152000
Deferred revenue—insurance 7500 =150000*7600/152000